OpenStudy (anonymous):

Find the area of a parallelogram if a base and corresponding altitude have the indicated lengths. Base 1 1/2 feet, altitude 6 inches. Area = sq. ft.

OpenStudy (a_clan):

area = base * height

OpenStudy (anonymous):

altitude = 6/12 ft area = base*height = 3/2 * 6/12 sqft = 18/24 ft = 3/4 sqft

OpenStudy (candyme):

Because Area=Base*Height, you would set this Equation up: A=18*6 (This is all inches) A=108 Inches But... you need feet, so 108/12= 9 so this is 9 square feet.

OpenStudy (anonymous):

@candyme u r making a very basic mistake.... first of all area = 108 sq.inch soooo, u hv to convert it to sqft to do that u hv to divide by 12*12 108 3 so Area in sq ft = ------- = ---- !!!!!! ☺ 12*12 4

OpenStudy (anonymous):

so if u want the answer in a decimal form then it is 0.75 sqft